Without referring to the preprogrammed function on your financial calculator​, use the basic formula for present​ value, along with the given discount​ rate, r​, and the number of​ periods, n​, to calculate the present value of​ $1 in the case shown in the following table

r=17%, n=6
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1=P*(1.17)^6

P=1/(1.17)^6

=$0.39(Approx).